zarkam21:

Fill out the tables for each scenario and answer the question that follows. Use $7.25 as the minimum wage and remember that employees in the United States must be paid time-and-a-half (1.5 times the normal hourly rate) for each hour worked over 40 hours per week.

zarkam21:

1 attachment
zarkam21:

@Vocaloid

Vocaloid:

well I think the table is supposed to be filled out for the Mongolian workers right? so if three laborers are each paid $1.10 per day (yikes) and they work for six days a week 50 weeks a year how much do they get paid per year?

zarkam21:

330

Vocaloid:

good, that's the cost per person so the total cost of labor = $990 materials: 50 per rug, 12 rugs --> total materials cost = ?

zarkam21:

600

Vocaloid:

awesome and the fixed costs are already given as $10,000 so that's column 3 total revenue = (2000 dollars per rug) * 12 rugs = ?

zarkam21:

24000

Vocaloid:

good so total revenue = $24,000 profit = total revenue - fixed cost - materials cost - labor = ?

zarkam21:

12,410

Vocaloid:

good, now let's start working on the US workers the mongolian workers work 14 hours a day 6 hours a week = 84 hours per week which means 84 - 40 = 44 hours of overtime so can you try calculating weekly wage for a US worker who works 40 hours a week at minimum wage + 44 hours of overtime?

zarkam21:

40*1.10

zarkam21:

44*1.10

zarkam21:

44+48.4=92.4

Vocaloid:

oh hold on just a sec: $1.10 is the daily wage for the mongolian workers not US workers, we get paid more than that haha

zarkam21:

lol

Vocaloid:

the problem says to use $7.25 as the hourly minimum wage so start with the normal hours, 40 hours a week at minimum wage gives how much income?

zarkam21:

290

Vocaloid:

awesome now for the overtime, we get paid 1.5 * hourly wage for any hours over 44 so 44 hours of overtime gives how much income?

zarkam21:

319

Vocaloid:

hm 1.5*7.25*44 = ?

zarkam21:

478.5

Vocaloid:

good, so the total weekly wage for one person = 478.5 + 319 = 797.5 per week now we just multiply this by 3 workers * number of weeks per year (which they give as 50) so 797.5 * 3 * 50 = $119,625 as the total cost of labor

Vocaloid:

let me just double check those calculations real quick

zarkam21:

Sure.

Vocaloid:

oh whoops I added the wrong overtime # total weekly wage = (40)(7.25)+44*1.5*7.25 = 768.5 weekly wage * 3 workers * 50 weeks per year = $115,275 so $115,275 should be it

zarkam21:

1 attachment
zarkam21:

same thing riht

Vocaloid:

yeah, just be careful since we're not talking about the same time period (this one is over 1 week instead of 1 year) so if each of the five workers is paid $2.50 per day to work 6 days per week what's the total cost of labor

zarkam21:

5*2.50*6=75

Vocaloid:

good then for the second column the cost per shirt is $75/500 shirts = $0.15 per shirt

Vocaloid:

total revenue column = 500 shirts * 40 dollars per shirt = ?

zarkam21:

20000

Vocaloid:

good fixed cost = already given total profit = total revenue - fixed cost - total labor = ? (ignore the cost per shirt for this calculation)

zarkam21:

9,999.85

Vocaloid:

0.15 is the cost per shirt, the total cost of labor is 75 so 20000-10000-75 = $9,925 simply divide this by 500 shirts to get the profit per shirt (last column)

zarkam21:

19.85

zarkam21:

the column before the last is $9,925?

Vocaloid:

yes and yes

zarkam21:

oh wait the last part at the bottom

zarkam21:

7.75*5

Vocaloid:

well let's start by finding out how many hours they work per week 10 hours a day * 6 days per week = 60 hours per week so that gives us 40 normal hours and 20 overtime hours can you try calculating the total income? for normal hours the wage is $7.25 and for overtime hours the wage is 1.5 * 7.25

zarkam21:

40*7.25=290 20*1.5*7.25=217.5

Vocaloid:

A+ now we add those together to get the weekly wage per person 290 + 217.5 = 507.5 we have five workers so we quintuple this to get 5 * 507.5 = $2,537.5 as the total cost of labor